Cast Iron Grills, Inc., manufactures premium gas barbecue grills. The company uses a periodic inventory system and the LIFO cost method for its grill inventory. Cast Iron's December 31, 2018, fiscal year-end inventory consisted of the following (listed in chronological order of acquisition):

Units

Unit Cost

5,600

$

400

4,300

500

6,600

600


The replacement cost of the grills throughout 2019 was $700. Cast Iron sold 30,000 grills during 2019. The company's selling price is set at 200% of the current replacement cost.

Required:
1. & 2. Compute the gross profit (sales minus cost of goods sold) and the gross profit ratio for 2019 under two different assumptions. First, that Cast Iron purchased 31,000 units and, second, that Cast Iron purchased 16,500 units during the year.
4. Compute the gross profit (sales minus cost of goods sold) and the gross profit ratio for 2019 assuming that Cast Iron purchased 31,000 units (as per the first assumption) and 16,500 units (as per the second assumption) during the year and uses the FIFO inventory cost method rather than the LIFO method.
Answer #1

Ansoor Vid selling price = 700 x 200 $1400 LIFO method ir Assumption, sales = 30,000 x 1400 – $42000000 Cost di goods sold =Assumption , Sales 30,000 X 1400 = $ 42000000 cost as goods sold = 16,500 4700)4(6,600x600)+(2,300 x 500) +(1400x400) =(11550☺ FIFO prethods Assumption , sales = 30,000 X1U00 = $42000000 cost di goods sold = (5 600x400) + (4200 X500)+(6600 X 606) +(= 17800000 42000000 -17.800000 = $auaooooo Gross profil 4. - _24200000 xeoo 42000000 = 57.62%